OpenStudy (anonymous):

Commutative Property 1 : (3 + 8) + 4 = 3 + (8 + 4) : Multiplication Property of Zero 2 : 1(6) = 6 : Associative Property 3 : 5(3 + x) = 15 + 5x : Distributive Property 4 : 5(0) = 0 : Multiplicative Identity 5 : 5 + 2 = 2 + 5

OpenStudy (vishweshshrimali5):

YOU HAVE TO MATCH THEM . RIGHT? Commutative property : \[5+2 = 2+5\] Multiplication property of zero ; \[5(0) = 0\] Associative property : \[(3+8) + 4 = 3+(8+4)\] Distributive property : \[5(3+x) = 15+3x\] Multiplicative identity : \[1(6) = 6\]

OpenStudy (vishweshshrimali5):

Hope you got that !!